# Individual or collective treatments: how to target antimicrobial use to limit the spread of Mannheimia haemolytica among beef cattle?

### v1

@@ Abstract
This study aimed at evaluating the impact of various treatment practices on the spread of BRD, specifically focusing on criteria for implementing collective treatments. Using a mechanistic stochastic model, we simulated the spread of Mannheimia haemolytica in a multi-pen fattening operation under sixteen different scenarios, considering pen composition, individual risk levels, and treatment strategies. Our findings suggest that an alternative criterion for collective treatments based on the speed of the disease spread, could reduce BRD incidence and AMU more effectively than conventional methods. This research highlights the importance of responsible treatment practices and the potential benefits of novel criteria for collective treatment strategies in improving animal health. Moreover, it emphasizes the need for transparency on the exposure to risk factors along the production chain.
@@ Introduction
The general overuse of antibiotics is of growing concern worldwide as it is thought to decrease treatment effectiveness ([laxminarayan_antibiotic_2013, coetzee_association_2019]) . Indeed, excessive and inappropriate use of antimicrobials can select for resistant bacterial strains, making it harder to treat infections in both animals and humans ([woolums_multidrug_2018, chen_scoping_2022]).
@@
Despite considerable progress made in the livestock sector to reduce antimicrobial use (AMU) ([bateman_evaluation_1990, watts_antimicrobial_2010]) the cattle production system is still heavily reliant on this method to control BRD ([oconnor_mixed_2013, abell_mixed_2017]). Individual treatment when animals show signs of disease is mostly used to care for affected animals, but collective treatments are also widely embraced to prevent BRD outbreaks and minimize economic losses ([nickell_metaphylactic_2010, ollivett_brd_2020]). The latter involves the administration of appropriate medications, such as antibiotics or anti-inflammatory drugs, to affected animals as well as the entire group to prevent additional cases ([lees_rational_2002, ives_use_2015]). This solution is motivated by poor sensitivity of current BRD field diagnostic methods, typical pathogenesis of BRD, and labor issues ([ives_use_2015]). As a result, a trade-off exists between antimicrobial use and bovine respiratory disease (BRD) occurrence. This trade-off highlights the need for responsible antimicrobial use practices, such as targeted collective treatment strategies. Yet, to this day, such strategies have not been proposed. Indeed, the relevant criteria and associated threshold to implement collective treatments are still questioned ([edwards_control_2010]).
@@
Modeling is used to better understand complex infection dynamics and compare scenarios and rank interventions ([ezanno_how_2020]). In the case of BRD, a model ranking treatment practices at pen level has been published ([picault_modelling_2022]). It simulates the circulation of an average pathogen in conditions mimicking contrasted farming contexts such as small versus large pens (considered as isolated from other pens) and low versus high risk of contracting BRD at pen level. The authors concluded on a bigger BRD circulation in large pens composed of high risk individuals and discussed which pens could get the highest benefit from collective treatments. A second model compares the impact of farm management practices on the spread of three causal pathogens of BRD in multi-pen farms featuring several individual risk levels of contracting BRD ([sorin-dupont_modeling_2023]). This study highlighted that sorting animals in pens in accordance with their individual risk level of developing BRD participates in limiting the spread of the most contagious pathogens. However, no model integrating novel criteria for triggering collective treatments has been proposed so far.
@@
This study aimed at evaluating the impact of treatment practices on the spread of a bovine respiratory pathogen in contrasted farming scenarios. More specifically, the objective was to evaluate criteria for implementing collective treatments in order to identify the best criterion. To that end, we used a mechanistic stochastic model of the spread of Mannheimia haemolytica in a multi-pen fattening operation. This choice of pathogen was motivated by the fact that M. haemolytica is commonly reported during BRD episodes in the field ([klima_characterization_2012, sudaryatma_bovine_2018]). Furthermore, its response to antimicrobials is well characterized, making it a good candidate for a pathogen specific model investigating treatment strategies ([noyes_mannheimia_2015]). We investigated sixteen contrasted scenarios reflecting farming practices in fattening operations, based on pen composition, individual risk level distribution, and individual vs. collective treatments based on conventional and original criteria. We compared scenario outputs regarding case occurrence, severity, and AMU.
@@
We modeled the spread of M. haemolytica in a multi-pen fattening farm and the treatment of associated BRD. We extended our stochastic and mechanistic individual-based model previously published ([sorin-dupont_modeling_2023]) for studying the circulation of specific pathogens involved in BRD, including M. haemolytica, within a fattening farm building, by including the possibility of collective treatments based on conventional and original criteria. Our model enabled pathogen transmission in a building containing several pens, i.e., groups of animals sharing a same closed indoor space. We worked under the assumption that an animal in a given pen was mainly exposed to infectious animals within its pen while also being exposed to the animals in the other pens in the building, albeit to a lesser extent. Our model monitored 6 individual characteristics. First, the model featured individual risk status of developing BRD, which referred to an individual qualitative information (low, medium, high) summarizing the level of stressors predisposing to BRD that the animal had been exposed to prior to its arrival. The model also featured hyperthermia (either due to BRD or unspecific), health status (susceptible, asymptomatic carrier, infectious), clinical status (asymptomatic, or with mild or severe clinical signs), detection status (detected or not), and treatment status (treated or not treated). Hereafter, we detail how treatments are triggered in this new version of the model. A graphic overview of the model and details on other (unchanged) processes and transitions can be found in SI1 and SI2.
@@
The treatment process represented a standard antimicrobial treatment protocol during the fattening period. Similarly to previous models, animals could be detected upon exhibition of clinical signs or hyperthermia. Severe clinical signs were detected with a sensitivity of 1, while the sensitivity for mild clinical signs detection was averaged to 0.5 ([kayser_evaluation_2019]). We assumed a clinical check-up at every time step (12 h, in agreement with farmers’ herd monitoring frequency ([terry_strategies_2021]). After detecting the first case through visual appraisal, all hyperthermic animals were identified using rectal temperature measured at the next feeding time, 12 hours later. ([picault_modelling_2022, sorin-dupont_modeling_2023]). Any animals detected as affected, either though visual detection or hyperthermia, received one antibiotic dose, assumed to be effective after duration $\tau_{T}$. Animals still exhibiting clinical signs after this duration were treated again (same duration), the number of treatments per episode being limited ($max_{T}$). Transitions from T to NT occurred in three cases: (1) the animal recovered after duration $\tau_{T}$ due to the treatment success with probability $p_{T}$; (2) the end of the infectious period occurred while under treatment but was not caused by it; (3) the animal still expressed clinical signs but was already treated maxT times, thus was not treated again (treatment failure).
@@
The second criterion (severity criterion) consisted in triggering a collective treatment based on a severity function. A collective treatment was triggered when this function reached a threshold, which selection is defined in SI3. The proposed function is a weighted sum of the number of animals detected as having mild or severe clinical signs. Mathematically, this translates to equation (1):
@@
with $D_{Mild}$ the number of animals detected as having mild clinical signs, and $D_{C}$ the number of animals detected as having severe clinical signs. The weight parameter $\alpha$ was smaller than 0.5 to allow an emphasis on severe cases. The process for selecting a relevant value for $\alpha$ is detailled in SI3.
The third criterion (slope criterion) considered to trigger a collective treatment was based on the speed of case appraisal. A collective treatment was triggered when this function reached a threshold, which selection is defined in SI3. The slope at time t in pen i represented the ratio of the cumulative incidence of detected BRD cases in pen i divided by the time since the first case appraisal in pen i. This translates to equation (2):
@@
where $cumulateD_{i}$ is the cumulative number of individuals detected as affected in pen $i$, $t_{0}$ being the starting time of an episode (i.e. the date of the first detected case), and $t$ the current time. In all these cases, a collective treatment can be performed only once per pen over the simulation time.
@@
We investigated sixteen scenarios of BRD treatment in a building of 200 young beef animals, distributed in 10 pens of 20 animals each, during 40 days, as BRD typically occurs in the first few weeks of fattening ([timsit_transmission_2013, assie_exposure_2009]). These scenarios were defined by one of the four treatment strategies, applied in 2 possible farm-scale proportions of individual risk status of developing BRD and two ways of assigning animals to pens. The four treatment strategies were: (1) individual treatment administered upon detection of clinical signs (Individual); (2) collective treatment triggered when the pen-scale cumulative incidence of BRD cases reached the standard threshold, in addition to individual treatment (Conventional); (3) collective treatment triggered when the weighted severity reached the defined threshold, in addition to individual treatment (Severity); and (4) collective treatment triggered when the slope function reached the defined threshold, in addition to individual treatment (Slope). The two possible individual risk level distributions in the building were: (1) 30%, 40%, and 30% of low, medium, and high risk individuals respectively (HR30); and (2) 10%, 0% and 90% of low, medium and high risk individuals respectively (HR90) ([amrine_evaluation_2019]). Since having a majority of low risk individuals in the building would not feature enough cases to trigger a collective treatment, it was thus left out of this study. Finally, animals could either be assigned into pens randomly (Random) or according to their individual risk level (Sorted).
@@
Each scenario had 200 stochastic replicates. At the building scale (i.e. grouping the 10 pens), we observed the distribution and the median of the cumulative incidence of BRD cases and the average time spent with severe clinical signs per affected animal (in hours) with regards to the total AMU in the building. We also computed the treatment False Discovery Rate (FDR) for each scenario. This measure corresponds to the ratio of false positives over predicted positives, i.e in our case the ratio between the number of doses given to healthy animals over the total AMU. Furthermore, we observed the delay between the detection of the first BRD case and the collective treatment implementation. We also observed the delay between the infection and the administration of the first dose of antimicrobial on an individual scale. Finally, for the sorted scenario, we counted the proportion of pens of each risk level that triggered a collective treatment. For each of these outputs, all applicable scenarios were compared. As we could neither assume the normality of the distributions or the homogeneity of variances across the scenarios for most of the outputs, our analysis was limited to trends as assumptions for statistical tests were not verified.

### v2

@@ Abstract
This study aimed at evaluating the impact of various strategies of collective treatments with antimicrobials on the spread of a BRD pathogen, Mannheimia haemolytica, specifically focusing on criteria for implementing collective treatments. Using a mechanistic stochastic model, we simulated the spread of M. haemolytica in a multi-pen fattening operation under sixteen different scenarios, considering pen composition, individual risk levels, and treatment strategies. Our findings suggest that an alternative criterion for collective treatments based on the speed of the disease spread, could reduce BRD incidence and AMU more effectively than conventional methods. This research highlights the importance of responsible collective treatment strategies and the potential benefits of novel criteria for collective treatment in improving animal health. Moreover, it emphasizes the need for transparency on the exposure to risk factors along the production chain.
@@ Introduction
Excessive and inappropriate use of antimicrobials can select resistant bacterial strains, making it harder to treat infections in both animals and humans ([woolums_multidrug_2018, chen_scoping_2022]). However, antimicrobial treatments remain the main control measure for health practioners to treat bacterial infectious diseases, thus making antimicrobioresistance a global concern ([laxminarayan_antibiotic_2013, coetzee_association_2019]).
@@
While viral pathogens are typically more infectious, much of the internal damage and clinical signs associated with BRD arise from secondary bacterial infections ([rice_mannheimia_2007]). Pasteurellaceae such as Mannheimia haemolytica or Pasteurella multocida are often associated with severe BRD cases, based upon necropsies of animals presenting severe respiratory symptomes, as well as positive serologies ([booker_microbiological_2008, welsh_isolation_2004]). M.haemolytica is particularly recurrent, with seroprevalences up to 80% and incidence risks up to 22.7% ([assie_exposure_2009]). M. haemolytica is commonly reported during BRD episodes in the field ([klima_characterization_2012, sudaryatma_bovine_2018]). Curative treatments for BRD mostly focus on these secondary infections, as they consist in using antimicrobials ([edwards_control_2010]). M.haemolitica’s response to antimicrobials is well characterized, making it a good candidate for a pathogen specific studies investigating treatment strategies ([noyes_mannheimia_2015]).
@@
Despite considerable progress made in the livestock sector to reduce antimicrobial use (AMU) ([bateman_evaluation_1990, watts_antimicrobial_2010]) the cattle production system is still heavily reliant on this method to control BRD ([oconnor_mixed_2013, abell_mixed_2017]). Pathogen identification is not done on routine, therefore antimicrobial curative treatment is the baseline response in the event of an outbreak ([dedonder_review_2015]). Individual treatment when animals show signs of disease is mostly used to care for affected animals, but collective treatments are also widely embraced to prevent BRD outbreaks and minimize economic losses ([nickell_metaphylactic_2010, ollivett_brd_2020, baggott2011demonstration]). The latter involves the administration of appropriate medications, such as antibiotics or anti-inflammatory drugs, to affected animals as well as the entire group to prevent additional cases ([lees_rational_2002, ives_use_2015]). This solution is motivated by poor sensitivity of current BRD field diagnostic methods, typical pathogenesis of BRD, and labor issues ([ives_use_2015]). Indeed, the sensitivity for BRD diagnosis was evaluated at around 0.27, making it very unefficient to treat each animal individually upon visual appraisal ([timsit2016diagnostic]). As a result, a trade-off exists between antimicrobial use and bovine respiratory disease (BRD) occurrence. This trade-off highlights the need for responsible antimicrobial use practices, such as targeted collective treatments with antimicrobials. Yet, to this day, such strategies have not been proposed. Indeed, the relevant criteria and associated threshold to implement collective treatments are still questioned ([edwards_control_2010]).
@@
Modeling is used to better understand complex infection dynamics and compare scenarios and rank interventions ([ezanno_how_2020]). In the case of BRD, a model ranking treatment strategies at pen level has been published ([picault_modelling_2022]). It simulates the circulation of an average pathogen in conditions mimicking contrasted farming contexts such as small versus large pens (considered as isolated from other pens) and low versus high risk of contracting BRD at pen level. The authors concluded on a bigger BRD circulation in large pens composed of high risk individuals and discussed which pens could get the highest benefit from collective treatments. A second model compares the impact of farm management practices on the spread of three causal pathogens of BRD in multi-pen farms featuring several individual risk levels of contracting BRD ([sorin-dupont_modeling_2023]). This study highlighted that sorting animals in pens in accordance with their individual risk level of developing BRD participates in limiting the spread of the most contagious pathogens. However, no model integrating novel criteria for triggering collective treatments has been proposed so far.
@@
This study aimed at evaluating the impact of collective treatment strategies on the spread of M.haemolytica in contrasted farming scenarios. More specifically, the objective was to evaluate criteria for implementing collective treatments in order to identify the best criterion. To that end, we used a mechanistic stochastic model of the spread of Mannheimia haemolytica in a multi-pen fattening operation. We investigated sixteen contrasted scenarios reflecting farming practices in fattening operations, based on pen composition, individual risk level distribution, and individual vs. collective treatments based on conventional and original criteria. We compared scenario outputs regarding case occurrence, severity, mortality and AMU.
@@
We modeled the spread of M. haemolytica in a multi-pen fattening farm and the treatment of associated BRD. We extended our stochastic and mechanistic individual-based model previously published ([sorin-dupont_modeling_2023]) for studying the circulation of specific pathogens involved in BRD, including M. haemolytica, within a fattening farm building, by including the possibility of collective treatments based on conventional and original criteria. Our model monitored 6 individual characteristics, each animal being therefore defined by its state in each of the 6 state machines included in the model. First, the model featured individual risk status of developing BRD, which referred to an individual qualitative information (low, medium, high) summarizing the level of stressors predisposing to BRD that the animal had been exposed to prior to its arrival. The model also featured hyperthermia (either due to BRD or unspecific), infection status (susceptible, asymptomatic carrier, infectious), clinical status (asymptomatic, or with mild or severe clinical signs), detection status (detected or not), and treatment status (treated or not treated). Hereafter, we detail how treatments are triggered in this new version of the model. A graphic overview of the model is proposed in Figure 1.
Figure 1: Graphic overview of the states and transitions in each process of the model. Each animal is defined by its state in each of the 6 state machines included in the model. Full arrows represent a transition from one state to another whereas the dashed lines represent regulatory links. Let $X$ be a regulator (state or transition) of transition $T_{2}$. This will be represented by a dashed arrow starting from $X$ and targeting $T_{2}$. from [sorin-dupont_modeling_2023]
Hyperthermia
Hyperthermia was composed of two states: hyperthermic (H) and non-hyperthermic (NH). NH animals could transition to H with probability $p_{H}$ attributed to non-infectious factors. Once in state H, they remained there for a period $\tau_{H}$ sampled from a Beta distribution adjusted using observed data, before reverting to NH. Additionally, the transitions from NH to H and back could result from the infection process with a 1 probability.
Infection and clinical signs status
Four infection statuses were considered: susceptible (S), asymptomatic carrier (E), infectious (I) and recovered (R) animals. Asymptomatic carriers were modeled as non infectious animals which carry M.haemolytica commensally. They could spontaneously turn I with probability $p_{E}$ and could also be infected by surrounding infectious individuals (I). Three actions were triggered when entering the state I: (1) the individual exhibited mild clinical signs for a duration $\tau_{M}$ drawn from a Beta distribution calibrated from observed data, (2) the animal changed from NH to H state, (3) a random draw with probability $p_{C}$ drove whether the individual would display severe clinical signs at the end of its mild clinical signs. The probability $p_{C}$ is higher if the animal has a high individual risk level. If displaying severe clinical signs, a boolean deciding on the survival of the individual was drawn from a binomial law of probability $p_{d}$. Death occurred at the end of the severe clinical signs duration ($\tau_{C}$). If the individual did not die and was still infectious after $\tau_{C}$, it transitioned back to MildC. The animal then recovered and became R at the end of the infectious period. Recovery occurred after a duration $\tau_{I}$ drawn from a gamma distribution. Theoretically, $\tau_{I}$ is longer than $\tau_{M}$. However, $\tau_{M}+\tau_{C}$ could exceed $\tau_{I}$. In that case, the infectious period was $\tau_{M}+\tau_{C}$. When transitioning to R state, animals changed from H to NH.
Detection status
@@
Similarly to previous models, animals could be detected upon exhibition of clinical signs or hyperthermia. Severe clinical signs were detected with a sensitivity of 0.6, while the sensitivity for mild clinical signs detection was averaged to 0.3, assuming that lethargy is the main driver for detection of severe clinical signs ([kayser_evaluation_2019]). We assumed a clinical check-up at every time step (12 h, in agreement with farmers’ herd monitoring frequency ([terry_strategies_2021]). After detecting the first case through visual appraisal, all hyperthermic animals were identified using rectal temperature measured at the next feeding time, 12 hours later. ([picault_modelling_2022, sorin-dupont_modeling_2023]). False discovery may occur with the temperature check as hyperthermia may be unrelated to infection.
Treatment status
The treatment process represented a standard antimicrobial treatment protocol during the fattening period. Each animal that was detected as affected, either though visual detection or hyperthermia, received one antibiotic dose, assumed to be effective after duration $\tau_{T}$ with a probability $p_{T}$, accounting for antimicrobial resistance. Animals still exhibiting clinical signs after this duration were treated again (same duration), the number of treatments per episode being limited ($max_{T}$). Transitions from T to NT occurred in three cases: (1) the animal recovered after duration $\tau_{T}$ due to the treatment success with probability $p_{T}$; (2) the end of the infectious period occurred while under treatment but was not caused by it; (3) the animal still expressed clinical signs but was already treated $max_{T}$ times, thus was not treated again and ignored from further treatments, as per on-field recommendations ([van2017reducing]). The values of the parameters as well as their source are summarized in Table 1.
Parameter Symbol Value Source
Transmission rate to susceptible (/h) $\beta$ 0.005, 0.008 and 0.012 for low, medium and high risk levels respectively [picault_modelling_2022]
Spontaneous shedding probability $p_{E}$ 0.14, 0.42 and 0.68 for low, medium and high risk levels respectively [frank_colonization_1986]
Initial carrier proportion $E_{0}/N_{0}$ 0.26, 0.48 and 0.72 for low, medium and high risk levels respectively [timsit_transmission_2013]
Mild clinical sign duration $\tau_{M}$ 2-8 (med=6) days [grissett_structured_2015]
Infectious period duration $\tau_{I}$ 6-10 (med=8) days [thomas_insights_2019]
Asymptomatic period duration $\tau_{E}$ 1 day [grissett_structured_2015]
Probability of successful treatment $p_{T}$ 0.71 [dedonder_review_2015]
Probability of severe forms $p_{C}$ 0.65 [timsit_transmission_2013]
Contact rate between pens $c$ 0.01 Assumed
Table 1: Parameter values used for M. haemolytica
Force of infection
Our model enabled pathogen transmission in a building containing several pens, i.e., groups of animals sharing a same closed indoor space. We assumed that an animal in a given pen was mainly exposed to infectious animals within its pen while also being exposed to the animals in the other pens in the building, albeit to a lesser extent.
The force of infection $\Phi_{i}$ for pen $i$ was frequency-dependent, taking into account the multi-pen nature of our model by accounting for the force of infection within the pen and between pens. Let $\mathbb{B}$ be the set of pens of $n$ individuals each raised simultaneously in a building. In each pen $i$, the intra pen force of infection was the sum of the individual contribution $\beta_{\rho}$ of each infectious set of individual of risk level $\rho$ ($I_{\rho,i}$). The contribution increased with the individual risk level $\rho$. Mathematically, this translates to equation (1). In a given pen $i$, the susceptible individuals experienced an intra pen force of infection with the addition of the contributions of the other pens of set $\mathbb{B}$ deprived of i ($\mathbb{B}\backslash\{i\}$) multiplied by a contact rate between pens $c$ defining how isolated the pens were from each other. In our simulations, we set $c=0.01$. This choice was motivated by the analysis of the effect of this parameter in the previous model ([sorin-dupont_modeling_2023]). This value implied a low contact rate between pens, which was compliant to the low airborne transmission of M.haemolytica ([timsit_transmission_2013]). A susceptibility factor $\sigma_{\rho}$ increasing with the individual risk level $\rho$ multiplied this total. Mathematically, this translates to equation (2). EMULSION automatically converts the force of infection $\Phi_{i}$ to a time-dependent probability $p_{\Phi_{i}}$ using the equation (3), with $\delta t$ being the current time step.
$\forall i\in\mathbb{B},\lambda_{i}=\sum_{\rho\in\mathbb{P}}{\beta_{\rho}I_{\rho,i}}$ (1)
$\Phi_{i}=\sigma_{\rho}\frac{\lambda_{i}+c\sum_{b\in\mathbb{B}\backslash\{i\}}{\lambda_{b}}}{N_{i}+c(N-N_{i})}$ (2)
$p_{\Phi_{i}}(\delta t)=e^{1-\Phi_{i}\delta t}$ (3)
with $\mathbb{P}$ the set of individual risk levels, $\beta_{\rho}$ the shedding level of individuals with risk level $\rho$, $I_{\rho,i}$ the number of infectious individuals with risk level $\rho$ in pen $i$, $N_{i}$ the total number of individuals in pen $i$, $N$ the total population size. This function was proposed for modeling the force of infection as it enables to represent perfectly separated pens ($c=0$) as well as a unique large pen with homogeneous contacts ($c=1$) using a single formula.
Collective treatment criteria
@@
The second criterion (severity criterion) consisted in triggering a collective treatment based on a severity function. A collective treatment was triggered when this function reached a threshold, which was selected in order to have the best trade-off possible between case incidence and antimicrobial use. The full selection process is defined in SI2. The proposed function is a weighted sum of the number of animals detected as having mild or severe clinical signs in a given pen $i$. Mathematically, this translates to equation (4):
@@
with $D_{Mild,i}$ the number of animals detected as having mild clinical signs in pen $i$, and $D_{C},i$ the number of animals detected as having severe clinical signs. The weight parameter $\alpha$ was smaller than 0.5 to allow an emphasis on severe cases. The process for selecting a relevant value for $\alpha$ is detailled in SI2.
The third criterion (slope criterion) considered to trigger a collective treatment was based on the speed of case appraisal. A collective treatment was triggered when this function reached a threshold, which selection answered to the same logic as the severity criterion. The slope at time $t$ in pen $i$ represented the ratio of the cumulative incidence of detected BRD cases in pen $i$ divided by the time since the beginning of the fattening. To avoid dividing by 0, $P_{i}(t)$ is only defined for $t>t_{0}$. This translates to equation (5):
@@
where $cumulateD_{i}$ is the cumulative number of individuals detected as affected in pen $i$, $t_{0,i}$ being the starting time of the simulation, and $t$ the current time. In all these cases, a collective treatment can be performed only once per pen over the simulation time.
@@
We investigated sixteen scenarios of BRD treatment in a building of 200 young beef animals, distributed in 10 pens of 20 animals each, during 40 days, as BRD typically occurs in the first few weeks of fattening ([timsit_transmission_2013, assie_exposure_2009]). These scenarios were defined by one of the four treatment strategies, applied in 2 possible farm-scale proportions of individual risk status of developing BRD and two ways of assigning animals to pens. The four treatment strategies were: (1) individual treatment administered upon detection of clinical signs (Individual); (2) collective treatment triggered when the pen-scale cumulative incidence of BRD cases reached the standard threshold, in addition to individual treatment (Conventional); (3) collective treatment triggered when the weighted severity reached the defined threshold, in addition to individual treatment (Severity); and (4) collective treatment triggered when the slope function reached the defined threshold, in addition to individual treatment (Slope). The two possible individual risk level distributions in the building were: (1) 30%, 40%, and 30% of low, medium, and high risk individuals respectively (HR30); and (2) 10%, 0% and 90% of low, medium and high risk individuals respectively (HR90). We chose these proportions in order to display contrasted situations, one being compliant with previously estimated risk proportion, and the other corresponding to a worst-case scenario ([amrine_evaluation_2019]). Since having a majority of low risk individuals in the building would not feature enough cases to trigger a collective treatment, it was thus left out of this study. Finally, animals could either be assigned into pens randomly (Random) or according to their individual risk level (Sorted).
@@
Each scenario had 200 stochastic replicates. At the building scale (i.e. grouping the 10 pens), we observed the distribution and the median of the cumulative incidence of BRD cases and the average time spent with severe clinical signs per affected animal (in hours) with regards to the total AMU in the building. Moreover, we compared the mortality accross the scenarios. We also computed the proportion of treatment misuse for each scenario. This measure corresponds to the ratio of false positives over predicted positives, i.e in our case the ratio between the number of doses given to healthy animals over the total AMU. Furthermore, we observed the delay between the detection of the first BRD case and the collective treatment implementation. We also observed the delay between the infection and the administration of the first dose of antimicrobial on an individual scale. Finally, for the sorted scenario, we counted the proportion of pens of each risk level that triggered a collective treatment. For each of these outputs, all applicable scenarios were compared. As we could neither assume the normality of the distributions or the homogeneity of variances across the scenarios for most of the outputs, non-parametric tests were used for the analysis. General effects of treatment modalities were assessed with Kruskal-Wallis tests and post-hoc pairwise comparisons were done by using Wilcoxon-Mann-Whitney with Bonferroni correction to reduce the family-wise error rate.
Sensitivity analysis
To better understand the behavior of the model and to characterize the impact of parameter uncertainty, we carried out an exploration of the model sensitivity by performing a one-at-a-time sensitivity analysis on six scenarios with a common risk level proportion of 30, 40 and 30% of low risk, medium risk and high risk respectively, as it was the most realistic ([amrine_evaluation_2019]). We simulated our three collective treatment strategies in randomly allocated pens and in sorted pens.
The model was derived from a previous one on which a sensitivity analysis accounting for parameter interaction was performed ([sorin-dupont_modeling_2023]). The latter did not provide evidence that parameter interactions contributed to model output variance. Therefore, our analysis was limited to a one-by-one analysis on the thresholds for our twonew collective treatment criteria as well as the two key parameters of disease transmission: the transmission rate $\beta$ and the contact rate between pens $c$. The thresholds and the transmission rates were used at their nominal values $\pm$ 10 and 25%. We used a larger interval for the contact rate, as little effect was observed in the sensitivity analysis of the previous model: we used the nominal value (0.01), multiplied or divided by 0.1, 0.5, 5 and 10. The tested values for each parameter are summarized in Table 2.
Parameter Nominal value Values
[…]
Table 2: Parameter values in the sensitivity analysis.
Each simulation had 200 stochastic replicates. We recorded the cumulative incidence and AMU. In order to evaluate the impat of a change in value of the transmission rate and the contact rate in each specific scenario, we computed the median of the simulations with the nominal values of the studied parameters. These medians were used as baselines. For each parameter change, we computed the relative variation of the output with regard to its respective baseline in order to evaluate the impact of the parameter change in value. We also compared the distributions of the outputs in the different collective treatment strategies with the tested parameter values in order to test the robustness of our previous conclusions. The distributions of the outputs were compared using Wilcoxon-Mann-Whitney tests with a Bonferroni correction, as normality could not be assumed.